OpenStudy (anonymous):

which equation is equivalent to 5x-2(7x+1)=14x...Pleasee Show Worky ;]

OpenStudy (anonymous):

which equation? 5x-14x-2 = 14x

OpenStudy (anonymous):

All i did was distribute the -2

OpenStudy (anonymous):

ightt' thank yhu :]

OpenStudy (anonymous):

its wrong tho cuz theres no answer for that

OpenStudy (anonymous):

5x-14x-2 = 14x -23x - 2 = 0? -23x = -2? 23x = 2? x = 2/23? these are all equivalent

OpenStudy (anonymous):

5x-14x-2 = 14x -9x - 2 = 14x -2 = 14x+9x -2 = 23x x =- 2/23?